OpenStudy (anonymous):

What would be the y-intercepts of the functions y=5^x and y=log5(x)?

OpenStudy (anonymous):

someone please help omg

OpenStudy (mathmale):

to find the y-intercept of a function like this, the general approach is to set x= to 0 and find y. Note that this is not true of the log function, but is true of y=5^x. Please evaluate y=5^x when x=0.

OpenStudy (alexandervonhumboldt2):

hint: a^0=1

OpenStudy (anonymous):

y=5^0 would be equal to 1

OpenStudy (anonymous):

So the y-intercept of y=5^x would be 1?

OpenStudy (mathmale):

In regard to \[y=\log_{5} x, \] why don't you graph this function and determine from the graph whether / where a y-intercept exists? To helpers: this question was aimed at @mysterylizard.

OpenStudy (mathmale):

Verify that the y-int. of 5^x is 1 by graphing it.

OpenStudy (mathmale):

Mystery?

OpenStudy (anonymous):

I graphed log5(x) and it kind of looks like it crosses the y-intercept at -1, but I can't tell 100%

OpenStudy (mathmale):

Have a graphing calculator? You can look up graphs of many different functions simply by searching the 'Net. Try searching for "log function."

OpenStudy (anonymous):

I don't have my graphing calculator with me, so I looked up the graph on the Internet. However, it didn't include exactly where the y-intercept is. I only guessed it was -1 from looking at the graph. Like I said, I'm not sure.

OpenStudy (mathmale):

Actually, Mystery, the log function never crosses the y-axis, and thus has no y-intercept. On the other hand, y=x^5 definitely has such an intercept; it is (0,1).

OpenStudy (anonymous):

Okay, thank you so much